The width of the garden is (x-5) feet and the length of the garden is 2 feet longer than the width. Write an expression for the area of the garden
Answer by Theo(13342)

w = width
l = length
w = x-5
l = x-5 + 2 = x - 3
area = l * w
area = (x-3) * (x-5)
when x = 5, the width is 0 which is not valid.
when x < 5, the width is negative which is not valid.
any value of x > 5 will be valid.
for example:
when x = 6, w = 1 and l = 3 and area = 3
when x = 7, w = 2 and l = 4 and area = 8
when x = 8, w = 3 and l = 5 and area = 15
etc.